OpenStudy (kenzkenz):

the midpoint of uv is (5, -11) the coordinates of one endpoint are u(3,5) find the coordinates of v

OpenStudy (sean_will21):

hey maybe i can help

OpenStudy (kenzkenz):

That would be great I'm really stuck on it

OpenStudy (sean_will21):

is that the full question or is there a way you can show me the full thing

OpenStudy (kenzkenz):

That's the full question

OpenStudy (sean_will21):

ohh okayy let me see

OpenStudy (sean_will21):

whats the answers for the question?

OpenStudy (kenzkenz):

There aren't any its a show your work problem

OpenStudy (sean_will21):

ok im doing the math rn

OpenStudy (sean_will21):

idk if this could be right and remember i said i could probably help and think its 2

OpenStudy (sshayer):



OpenStudy (sean_will21):

3x over 2 and then 5y over 2 thats all i can conclude from that

OpenStudy (sshayer):

3+x=10 x=10-3=7 5+y=-22 y=-22-5=?

OpenStudy (sean_will21):

hmmm...

OpenStudy (sean_will21):

ohh i see

OpenStudy (sean_will21):

y = -27

OpenStudy (sshayer):

so coordinates of v are (7,-27)

OpenStudy (sean_will21):

where did u get 7 from

OpenStudy (sshayer):

i have calculated above x=7

OpenStudy (sean_will21):

ohh yes the answer should be (7,-27)

OpenStudy (sean_will21):

u there?